# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ 4.2 Wilson fermions and reflection positivity
Fix Euclidean gamma matrices $\{\gamma_{\mu}\}_{\mu=1}^{d}$ with $\{\gamma_{\mu},\gamma_{\nu}\}=2\delta_{\mu\nu}$, and define $\Theta$ on the fermionic generators by
$(\Theta\psi)(n):=\gamma_{1}\,\bar{\psi}(\vartheta n)^{\sf T},\qquad(\Theta\bar{\psi})(n):=\psi(\vartheta n)^{\sf T}\gamma_{1},$ (4.17)
where T denotes transpose in the spin/internal indices (no complex conjugation is applied to the Grassmann generators themselves). On products and complex scalars we impose the standard graded anti-automorphism rules
@@
so $\Theta$ is anti-linear and reverses the order of Grassmann factors. With these conventions, $\Theta$ maps $\mathcal{A}_{+}$ into the corresponding algebra on $\Lambda_{-}$, and $\Theta^{2}=\mathrm{id}$ on $\mathcal{A}_{+}^{\mathrm{even}}$. As before, the dot in $\langle\Theta F\cdot F\rangle$ denotes the (pointwise) product in this graded algebra.
Let $\Delta=\{\pm 1,\dots,\pm d\}$ be the oriented direction set, and write $|\mu|\in\{1,\dots,d\}$ for the underlying coordinate direction. We use the conventions
$\gamma_{\mu}=\mathrm{sgn}(\mu)\,\gamma_{|\mu|},\qquad U_{-|\mu|}(n)=U_{|\mu|}(n-\widehat{|\mu|})^{-1}.$
Let $V(n;x)>0$ be a local scalar built from the embedding (e.g. a one-frame cell volume), let $M(n;x)$ be a local real mass/volume weight, and let $w_{\mu}(n;x)>0$ be a local link weight attached to the oriented link $(n,\mu)$, built from local Euclidean invariants of $x$. (Reflection positivity uses only locality and positivity; in applications one fixes engineering dimensions by choosing $w_{\mu}(n;x)\propto|e_{\mu}(n)|^{-1}$ up to dimensionless shape factors and absorbing the remaining normalization into $\kappa$.)
Fix Wilson parameter $r=1$ and define the prototype Wilson action [24] by
$S_{W}^{\circ}[x,U,\bar{\psi},\psi]$ $:=\sum_{n\in\Lambda_{T,L}}V(n;x)\,\bar{\psi}(n)\,M(n;x)\,\psi(n)$ (4.19)
[…]
Gauge invariance is term-by-term, so it suffices to check it for the prototype action.
Finally, average it with its $\theta$-partner
$S_{W}[x,U,\bar{\psi},\psi]:=\frac{1}{2}\Bigl(S_{W}^{\circ}[x,U,\bar{\psi},\psi]+S_{W}^{\theta}[x,U,\bar{\psi},\psi]\Bigr).$ (4.21)
By construction this action is local on the abstract lattice. Moreover, with $\Theta$ acting on $(x,U,\psi,\bar{\psi})$ by the standard OS rules), one has $\Theta S_{W}^{\circ}=S_{W}^{\theta}$, hence the averaged action $S_{W}$ is $\Theta$–invariant.
The map induced by $\Theta$ on the set of Grassmann generators is a site permutation composed with a fixed linear transformation on the spin/internal indices. In particular, it has unit Berezin Jacobian, and any reordering sign is irrelevant on $\mathcal{A}_{+}^{\mathrm{even}}$. Hence
@@
Let $S_{b}=S_{x}+S_{\mathrm{fields}}^{(\mathrm{bos})}$ denote the bosonic part of the action (geometry + gauge + bosonic matter) and let $S_{W}$ be the $\Theta$-averaged Wilson action (4.21). Set $S_{\mathrm{tot}}=S_{b}+S_{W}$ and assume that $S_{\mathrm{tot}}$ is local, $\Theta$-invariant, and has finite interaction range in the index-time direction. Then
$\langle\Theta F\cdot F\rangle\geq 0\qquad\text{for all }F\in\mathcal{A}_{+}^{\mathrm{even}},$
@@
$1_{\{x\in\mathcal{X}_{\mathrm{adm}}\}}e^{-S_{\mathrm{tot}}}\,Dx\,DU\,D\Phi\,D\bar{\psi}\,D\psi$
on $\Lambda_{T,L}$.
@@
The proof follows the same factorization through a finite reflection slab as in Theorem 4.3, using $\Theta$-invariance of the domain and product measures (including (4.22)) and the decomposition $S_{\mathrm{tot}}=S_{+}+S_{0}+S_{-}$ with $S_{-}=\Theta S_{+}$. The only additional input is the standard Wilson structure of the time-like nearest-neighbour couplings across the reflection plane: with $r=1$ these couplings occur with the orthogonal projectors
$P_{\pm}=\tfrac{1}{2}(1\pm\gamma_{1}),$
so the slab Berezin pairing contracts each even Grassmann monomial with its $\Theta$-conjugate with a nonnegative coefficient (using only positivity of the local link weights and unitarity of the time-like link variables). This yields $\langle\Theta F\cdot F\rangle\geq 0$ for all $F\in\mathcal{A}_{+}^{\mathrm{even}}$. ∎
@@ 4.3 Symanzik effective theory and universality
$Z(a)=\int_{\mathcal{X}_{\rm adm}}Dx\int DU\int D\Phi\;\exp\bigl(-S_{x}[x]-S_{\rm fields}[x,U,\Phi]\bigr).$ (4.23)
@@
The Symanzik effective field theory organizes long-distance correlators at fixed physical separation as an expansion in the cutoff $a$:
@@
where the operators $\mathcal{O}_{i}$ are local continuum operators compatible with the exact symmetries of the regulator, cf. [15, 16]. In particular, on a rigid hypercubic lattice the symmetry is only the hypercubic group, and this allows rotation-breaking operators (and additional operator mixing) that are forbidden by full $\mathrm{SO}(d)$ covariance.
@@
Hypothesis 4.6 (Short-range geometry of SR).
@@
This hypothesis is a genuine condition on the geometry measure and therefore constrains the admissible choices of $S_{x}$. A natural way to target (SR) is to include an explicit local stiffness scale for the embedding fluctuations (e.g. quadratic penalties around a reference metric as in (2.19)), which makes the non–zero–mode fluctuations of $x(n)=an+\xi(n)$ effectively massive in lattice units. The hard admissibility constraints are compatible with such a massive regime and help by excluding degenerate cells, but by themselves they do not guarantee a mass gap or exponential mixing. This distinction is particularly important in higher dimensions, where entropic instabilities of random–geometry ensembles are well known. Accordingly, we treat (SR) as an explicit assumption to be verified for any proposed $S_{x}$, and we include direct diagnostics in Section 5. In $d=2$ we find behaviour consistent with a short-range twisting regime; however the $d=4$ situation remains to be mapped.
Assuming (SR), one may integrate out the geometry field at fixed $a$ to obtain an effective action for the field sector $(U,\Phi)$ alone. Locality and (SR) imply that this effective action is local and admits a Symanzik expansion of the form (4.24). Moreover, because the DLR measure (on the quotient by global $\mathrm{SE}(d)$) is $\mathrm{SE}(d)$-covariant, the induced effective action for $(U,\Phi)$ is $\mathrm{SO}(d)$-covariant (and gauge invariant when gauge fields are present); in particular, only $\mathrm{SO}(d)$-scalar operators contribute to scalar observables. This removes the representation-theoretic source of rotation-breaking counterterms and, more importantly, potentially simplifies Symanzik improvement and operator renormalisation: the basis of allowed counterterms (and hence the pattern of operator mixing) is constrained by full $\mathrm{SO}(d)$ rather than only the hypercubic group.
@@
By contrast, DLR integrates over a local embedding field $x(n)$ that enters the action through local metric/volume factors. After quotienting by (or fixing) the global $\mathrm{SE}(d)$ zero modes to remove the infinite group volume—a global gauge choice affecting local correlators only through finite-volume effects—the geometry still fluctuates locally. Assuming (SR), integrating out $x$ produces a local effective action for $(U,\Phi)$ to which the Symanzik expansion applies, and the underlying $\mathrm{SE}(d)$ covariance implies that the induced effective action is globally $\mathrm{SO}(d)$-covariant (and gauge invariant when gauge fields are present).
A purely global ‘geometry’ variable (such as a rigid rotation) has infinite correlation length; if it were coupled nontrivially it would act as a mediator producing nonlocal effective interactions upon elimination, which lies outside the scope of the Symanzik expansion and is exactly what (SR) excludes.
To connect the Symanzik discussion to the usual perturbative intuition, write locally
$x(n)=a\,n+\xi(n),\qquad e_{\mu}(n)=x(n+\hat{\mu})-x(n)=a\,\hat{\mu}+\delta e_{\mu}(n),$ (4.25)
where $\delta e_{\mu}(n)=\xi(n+\hat{\mu})-\xi(n)$. Introduce the dimensionless first differences
$\varsigma_{\mu}\xi(n):=\frac{\xi(n+\hat{\mu})-\xi(n)}{a}=\frac{\delta e_{\mu}(n)}{a}\in\mathbb{R}^{d}.$ (4.26)
All local geometric data entering the field actions are smooth local functions of the finite collection $\{\varsigma_{\mu}\xi(n)\}_{\mu}$, more precisely, of the frame matrix $E(n)$ built from the $e_{\mu}(n)$. Expanding around the regular embedding therefore yields schematic local expansions
$g_{\mu\nu}(n)=a^{2}\Big(\delta_{\mu\nu}+O(\varsigma\xi)\Big),\qquad V(n)=a^{d}\Big(1+O(\varsigma\xi)+O((\varsigma\xi)^{2})\Big),$ (4.27)
with coefficients that are bounded uniformly on $\mathcal{X}_{\mathrm{adm}}$ by admissibility (uniform shape regularity).
Let us consider the scalar sector for definiteness. Writing the scalar action on a general admissible geometry as $S_{\phi}[x,\phi]=S_{\phi}^{(0)}[\phi]+S^{(1)}_{\phi\xi}[\phi,\xi]+S^{(2)}_{\phi\xi}[\phi,\xi]+\cdots$, the regular-lattice part $S_{\phi}^{(0)}$ is the usual hypercubic action, while the geometry-dependent pieces split into two distinct families already at first order, as follows.
@@
$S^{(1)}_{\phi\xi}\ \supset\ \sum_{n}a^{d}\;\Big(\varsigma\xi\Big)(n)\,(\nabla^{(0)}\phi)(n)\,(\nabla^{(0)}\phi)(n),\qquad S^{(2)}_{\phi\xi}\ \supset\ \sum_{n}a^{d}\;\Big(\varsigma\xi\Big)^{2}(n)\,(\nabla^{(0)}\phi)^{2}(n),$ (4.28)

### v2

@@ 4.2 Fermions and reflection positivity
Fix constant Euclidean gamma matrices $\{\gamma_{a}\}_{a=1}^{d}$ with $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}$. As in the standard lattice reflection-positivity setup, we take a representation in which the relevant transpose identities hold; concretely, it suffices that the gamma matrices are chosen so that the matrix involutions introduced below satisfy $\Xi(n)^{\sf T}=\Xi(n)$ and $\Xi(n)^{2}=I$.
Recall that the local frame matrix $E(n)$ is built from the outgoing edge vectors at $n$ and that $E^{\theta}(n)$ denotes the time-reflected corner frame (same spatial columns, but using the backward time edge). From a frame $E^{\sharp}(n)$, which may be either $E$ or $E^{\theta}$, we form the corresponding geometry-dependent Dirac matrices
$\Gamma_{\mu}^{\sharp}(n)=\sum_{a=1}^{d}\gamma_{a}\,(E^{\sharp}(n)^{-1})_{a\mu},\qquad\{\Gamma_{\mu}^{\sharp}(n),\Gamma_{\nu}^{\sharp}(n)\}=2\,g_{\sharp}^{\mu\nu}(n),$
so in particular $(\Gamma_{1}^{\sharp}(n))^{2}=g_{\sharp}^{11}(n)\,I$ and the normalized matrix $\widehat{\Gamma}_{1}^{\sharp}(n)={\Gamma_{1}^{\sharp}(n)}/{\sqrt{g_{\sharp}^{11}(n)}}$ is an involution, $(\widehat{\Gamma}_{1}^{\sharp}(n))^{2}=I$.
For the Osterwalder–Schrader reflection across the plane between $n_{1}=0$ and $n_{1}=1$ it is convenient to use, at each site, the normalized “time” matrix pointing toward the reflection plane:
$\Xi(n)=\begin{cases}\widehat{\Gamma}_{1}(n),&n\in\Lambda_{-}\ \textrm{with}\ n_{1}\leq 0,\\[3.00003pt] \widehat{\Gamma}_{1}^{\theta}(n),&n\in\Lambda_{+}\ \textrm{with}\ n_{1}\geq 1.\end{cases}$ (4.17)
This choice is entirely local (it uses only the corner frame attached to the site $n$) and is tailored so that the same geometric “time direction toward the plane” is used on both sides. Other local choices of $\Xi$ are possible; (4.17) is a particularly transparent one because it is an involution by construction and reduces to the constant matrix $\gamma_{1}$ on a regular embedding.
Now define $\Theta$ on the fermionic generators by
$(\Theta\psi)(n):=\Xi(n)\,\bar{\psi}(\theta n)^{\sf T},\qquad(\Theta\bar{\psi})(n):=\psi(\theta n)^{\sf T}\,\Xi(n),$ (4.18)
where T denotes transpose in the spin/internal indices (no complex conjugation is applied to the Grassmann generators themselves). On products and complex scalars we impose the graded anti-automorphism rules
@@
so $\Theta$ is anti-linear and reverses the order of Grassmann factors. With the involution properties of $\Xi$ (in particular $\Xi^{\sf T}=\Xi$ and $\Xi^{2}=I$), one has $\Theta^{2}=\mathrm{id}$ on $\mathcal{A}_{+}^{\mathrm{even}}$. As before, the dot in $\langle\Theta F\cdot F\rangle$ denotes the pointwise product in this graded algebra.
The map induced by $\Theta$ on the set of Grassmann generators is a site permutation composed with an invertible linear transformation on the spin/internal indices with determinant $\pm 1$. In particular, it has unit Berezin Jacobian, and any global reordering sign is irrelevant on $\mathcal{A}_{+}^{\mathrm{even}}$. Hence we have
@@
As in §3.1, we build fermion actions from local geometric data and finite stencils on the abstract lattice, and then apply the $\theta$–partner/averaging prescription to enforce $\Theta$–invariance. In particular, the averaged naive action (3.23) is local and has time-range one in the index–time direction.
As in the proof of Theorem 4.3, we decompose the extended field multiplet $\omega=(x,U,\phi,\bar{\psi},\psi)$ as $\omega=(\omega_{-},\omega_{0},\omega_{+})$ using a reflection slab $\Sigma_{\ell}$ wide enough to contain the time-range $\ell$ of the stencil. By finite time-range, no local term can simultaneously involve strict-past variables and strict-future variables; consequently one can (uniquely, up to regrouping purely $\omega_{0}$ terms) write
$S_{f}(\omega)=S_{f,+}(\omega_{0},\omega_{+})+S_{f,0}(\omega_{0})+S_{f,-}(\omega_{0},\omega_{-}),\qquad S_{f,-}=\Theta S_{f,+}.$ (4.21)
For fermions, locality and $\Theta$–invariance are not by themselves sufficient: the additional input is positivity of the slab pairing across the reflection plane, which we isolate next.
We say that a Grassmann-even fermion action $S_{f}$ has a reflection-positive slab form (of width $\ell$) if the slab weight admits a finite “sum of $\Theta$–squares” representation
$e^{-S_{f,0}(\omega_{0})}=\sum_{j\in J}(\Theta B_{j})(\omega_{0})\,B_{j}(\omega_{0}),\qquad B_{j}\in\mathcal{A}_{+}^{\mathrm{even}}\ \text{supported in }\Sigma_{\ell}\cap\Lambda_{+},$ (4.22)
for some finite index set $J$. (Finiteness is automatic since the Grassmann algebra on $\Sigma_{\ell}$ is finite-dimensional and the exponential truncates.) Equivalently, it suffices to have a representation $S_{f,0}=S_{f,0}^{\mathrm{loc}}-\sum_{k}c_{k}\,(\Theta\chi_{k})\chi_{k}$ with $c_{k}\geq 0$ bosonic and $\chi_{k}$ odd supported in $\Sigma_{\ell}\cap\Lambda_{+}$; expanding the exponential then yields (4.22).
@@
Let $S_{b}=S_{x}+S_{\mathrm{fields}}^{(\mathrm{bos})}$ be the bosonic part of the action and let $S_{f}$ be a Grassmann-even fermion action. Assume:
  • (i)
    $S_{b}$ and $S_{f}$ are local on the abstract lattice and have finite interaction range $\ell<\infty$ in the index-time direction.
  • (ii)
    $S_{b}(\Theta\omega)=S_{b}(\omega)$ and $S_{f}(\Theta\omega)=S_{f}(\omega)$, and the hard constraint $\mathcal{X}_{\mathrm{adm}}$ and the product measures are $\Theta$-invariant (including (4.20)).
  • (iii)
    $S_{f}$ admits a reflection-positive slab form of width $\ell$ in the sense of (4.21)–(4.22).
Then for all $F\in\mathcal{A}_{+}^{\mathrm{even}}$ one has
$\langle\,\Theta F\cdot F\,\rangle\ \geq\ 0,$
@@
$1_{\{x\in\mathcal{X}_{\mathrm{adm}}\}}e^{-S_{b}-S_{f}}\,Dx\,DU\,D\phi\,D\bar{\psi}\,D\psi.$
@@
As in Theorem 4.3, the finite time-range assumption implies a decomposition
$S_{b}(\omega)=S_{{b},+}(\omega_{0},\omega_{+})+S_{{b},0}(\omega_{0})+S_{{b},-}(\omega_{0},\omega_{-}),\qquad S_{{b},-}=\Theta S_{{b},+}.$
Combine this with (4.21) to obtain for the total action
$S(\omega)=S_{+}(\omega_{0},\omega_{+})+S_{0}(\omega_{0})+S_{-}(\omega_{0},\omega_{-}),\qquad S_{-}=\Theta S_{+},$
where $S_{0}=S_{{b},0}+S_{f,0}$ depends only on $\omega_{0}$.
Let $F\in\mathcal{A}^{\mathrm{even}}_{+}$, so $F$ depends only on $(\omega_{0},\omega_{+})$. Define the (graded) conditional functional
$\Psi_{F}(\omega_{0})=\int e^{-S_{+}(\omega_{0},\omega_{+})}\,F(\omega_{0},\omega_{+})\,d\omega_{+},$
where $d\omega_{+}$ includes the bosonic product measures and the Berezin factors over the $\omega_{+}$ Grassmann generators. Using $\Theta$-invariance of the domain and product measures and the relation $S_{-}=\Theta S_{+}$, one obtains the standard factorization
$\langle\Theta F\cdot F\rangle=\frac{1}{Z}\int e^{-S_{0}(\omega_{0})}\,(\Theta\Psi_{F})(\omega_{0})\,\Psi_{F}(\omega_{0})\,d\omega_{0}.$
Now split $e^{-S_{0}}=e^{-S_{{b},0}}\cdot e^{-S_{f,0}}$ and use (4.22):
$e^{-S_{f,0}}=\sum_{j\in J}(\Theta B_{j})B_{j}.$
Since $\Psi_{F}$ is even and $\Theta$ is an involution on the even algebra, the integrand becomes a finite sum of terms of the form
$e^{-S_{{b},0}(\omega_{0})}\;(\Theta(B_{j}\Psi_{F}))(\omega_{0})\,(B_{j}\Psi_{F})(\omega_{0}),$
each of which integrates to a nonnegative number by the defining property of a $\Theta$–square pairing on $\mathcal{A}_{+}^{\mathrm{even}}$. Summing over $j$ yields $\langle\Theta F\cdot F\rangle\geq 0$. ∎
Example 4.6 (Geometric Wilson/projector mechanism).
A standard sufficient mechanism for the slab-positivity condition (4.22) is that the couplings crossing the reflection plane factor through orthogonal projectors with nonnegative coefficients. In the present geometric setting, the natural constant projectors $P_{\pm}=\tfrac{1}{2}(1\pm\gamma_{1})$ are replaced by the site-wise projectors built from (4.17):
$P_{\pm}(n):=\tfrac{1}{2}\bigl(I\pm\Xi(n)\bigr),\qquad P_{\pm}(n)^{2}=P_{\pm}(n),\quad P_{+}(n)P_{-}(n)=0.$
Concretely, let $V(n;x)>0$ be a local scalar weight (e.g. a cell volume), let $M(n;x)\in\mathbb{R}$ be an ultralocal mass weight, and let $w_{\mu}(n;x)>0$ be a local link weight attached to the oriented link $(n,\mu)$, all built from local Euclidean invariants of $x$. Consider a Wilson-type prototype action whose time-like hops are written directly in projector form:
$S_{W}^{\circ}[x,U,\bar{\psi},\psi]$ $=\sum_{n\in\Lambda}V(n;x)\,\bar{\psi}(n)\,M(n;x)\,\psi(n)$ (4.23)
[…]
$\quad+(\text{terms not crossing the reflection plane}).$
Here $U_{-1}(n)=U_{1}(n-\hat{1})^{-1}$ as usual. The omitted terms may include arbitrary local spatial hops and on-slice couplings; they do not affect the slab pairing because they do not connect the two sides of the reflection plane.
Apply the $\theta$–partner/averaging prescription of §3.1 to obtain the averaged action $S_{W}=\tfrac{1}{2}(S_{W}^{\circ}+S_{W}^{\theta})$, hence $\Theta S_{W}=S_{W}$. In the slab decomposition (4.21), the only fermionic terms that genuinely couple the two sides are the time-like hops across the reflected time link, and because they factor through $P_{\pm}(n)$ with nonnegative coefficients, the corresponding slab weight admits a representation (4.22). Therefore $S_{W}$ satisfies hypothesis (iii) of Theorem 4.5, and reflection positivity holds for all $F\in\mathcal{A}_{+}^{\mathrm{even}}$.
@@ 4.3 Symanzik effective theory and universality
$Z(a)=\int_{\mathcal{X}_{\mathrm{adm}}(a)}Dx\int DU\int D\Phi\;\exp\bigl(-S_{x}[x]-S_{\rm fields}[x,U,\Phi]\bigr).$ (4.26)
@@
Symanzik effective field theory is a convenient way to parameterize lattice artifacts in long–distance observables, cf. [16, 17]. Concretely, one considers correlation functions of local operators at a fixed physical separation $|x|$ while taking the cutoff $a\to 0$. On the lattice this means that the separation in lattice units grows as $r=|x|/a\to\infty$, so the correlators probe distances that are large compared to the cutoff but still finite in physical units.
Under the standard assumptions of locality and a mass gap at the cutoff scale, the effect of the regulator at such distances can be encoded by a continuum effective action
@@
where $\{\mathcal{O}_{i}\}$ is a basis of local continuum operators (composite operators built from the fields and their derivatives) and $\Delta_{i}>0$ are their excess dimensions relative to the target action. The coefficients $c_{i}(g_{\rm bare})$ are matching coefficients: they depend on the microscopic definition of the regulator (hence on the bare couplings $g_{\rm bare}$) and are generally scheme–dependent, but the operator basis is not arbitrary. Indeed, the allowed operators $\mathcal{O}_{i}$ are precisely those compatible with the exact symmetries of the regulator.
In particular, on a rigid hypercubic lattice the exact symmetry is only the hypercubic group $\mathrm{H}(d)$, so the Symanzik operator basis is larger than the $\mathrm{SO}(d)$–scalar basis and radiative corrections can generate rotation–breaking higher–derivative terms. For a scalar field, writing $H_{\mu\nu}=\partial_{\mu}\partial_{\nu}\phi$, the $\mathrm{SO}(d)$–invariants at four-derivative order are $\mathrm{tr}(H)^{2}=(\partial^{2}\phi)^{2}$ and $\mathrm{tr}(H^{2})=\sum_{\mu,\nu}(\partial_{\mu}\partial_{\nu}\phi)^{2}$. By contrast, hypercubic symmetry permits independent coefficients for the diagonal and off–diagonal pieces, e.g., $\sum_{\mu}(\partial_{\mu}^{2}\phi)^{2}$ and $\sum_{\mu<\nu}(\partial_{\mu}\partial_{\nu}\phi)^{2}$. Reducing $\mathrm{SO}(d)$ to $\mathrm{H}(d)$ enlarges the allowed mixing among operators with the same quantum numbers, permitting additional anisotropic counterterms that are absent for an $\mathrm{SO}(d)$–covariant regulator.
@@
Hypothesis 4.7 (Short-range geometry).
@@
This hypothesis is a genuine condition on the geometry measure and therefore constrains the admissible choices of $S_{x}$. A natural way to target (SR) is to include an explicit local stiffness scale for the embedding fluctuations (e.g. quadratic penalties around a reference metric as in (2.21)), which makes the non–zero–mode fluctuations of $x(n)=an+\eta(n)$ effectively massive in lattice units. The hard admissibility constraints are compatible with such a massive regime and help by excluding degenerate cells, but by themselves they do not guarantee a mass gap or exponential mixing. This distinction is particularly important in higher dimensions, where entropic instabilities of random geometry ensembles are well known. Accordingly, we treat (SR) as an explicit assumption to be verified for any proposed $S_{x}$, and we include direct diagnostics in Section 5. In $d=2$ we find behaviour consistent with a short-range twisting regime; however the $d=4$ situation remains to be mapped.
Assuming (SR), one may integrate out the geometry field at fixed $a$ to obtain an effective action for the field sector $(U,\Phi)$ alone. Locality and (SR) imply that this effective action is local and admits a Symanzik expansion of the form (4.27). Moreover, because the DLR measure is $\mathrm{SE}(d)$-covariant, the induced effective action for $(U,\Phi)$ is $\mathrm{SO}(d)$-covariant, and gauge invariant when gauge fields are present. In particular, only $\mathrm{SO}(d)$-scalar operators contribute to scalar observables. This removes the representation-theoretic source of rotation-breaking counterterms and, more importantly, potentially simplifies Symanzik improvement and operator renormalisation: the basis of allowed counterterms (and hence the pattern of operator mixing) is constrained by full $\mathrm{SO}(d)$ rather than only the hypercubic group.
@@
By contrast, DLR integrates over a local embedding field $x(n)$ that enters the action through local metric/volume factors. After quotienting by (or fixing) the global $\mathrm{SE}(d)$ zero modes to remove the infinite group volume, the geometry still fluctuates locally. Assuming (SR), integrating out $x$ produces a local effective action for $(U,\Phi)$ to which the Symanzik expansion applies, and the underlying $\mathrm{SE}(d)$ covariance implies that the induced effective action is globally $\mathrm{SO}(d)$-covariant.
A purely global “geometry” variable (such as a rigid rotation) has infinite correlation length; if it were coupled nontrivially it would act as a mediator producing nonlocal effective interactions upon elimination, which lies outside the scope of the Symanzik expansion and is exactly what (SR) excludes.
To connect the Symanzik discussion to the usual perturbative intuition, expand the embedding about the regular one,
$x(n)=a\,n+\eta(n),\qquad e_{\mu}(n)=x(n+\hat{\mu})-x(n)=a\,\hat{\mu}+\delta e_{\mu}(n),$ (4.28)
where $\delta e_{\mu}(n)=\eta(n+\hat{\mu})-\eta(n)$ and $\eta(n)\in\mathbb{R}^{d}$. Throughout this and the next subsections, $\nabla_{\mu}$ denotes the forward lattice derivative
$(\nabla_{\mu}f)(n)=\frac{f(n+\hat{\mu})-f(n)}{a},$ (4.29)
acting componentwise on $\mathbb{R}^{d}$–valued fields. In particular,
$(\nabla_{\mu}\eta)(n)=\frac{\eta(n+\hat{\mu})-\eta(n)}{a}=\frac{\delta e_{\mu}(n)}{a}\in\mathbb{R}^{d}.$ (4.30)
Heuristically, if $\eta(n)=u(an)$ for a smooth interpolant $u$, then $\nabla_{\mu}\eta(n)=\partial_{\mu}u(an)+O(a)$.
Fix components with respect to the regular orthonormal frame $\{\hat{\mu}\}$:
$\eta_{\nu}(n)=\hat{\nu}\cdot\eta(n),\qquad(\nabla\eta)_{\mu\nu}(n)=\nabla_{\mu}\eta_{\nu}(n).$ (4.31)
Thus $\nabla\eta$ is the lattice Jacobian, its trace
$\nabla\!\cdot\eta(n)=\mathrm{tr}(\nabla\eta)(n)=\sum_{\mu}\nabla_{\mu}\eta_{\mu}(n),$ (4.32)
is the discrete divergence, and we write $|\nabla\eta|^{2}=\sum_{\mu,\nu}(\nabla_{\mu}\eta_{\nu})^{2}$ for the canonical $\mathrm{SO}(d)$–scalar quadratic in first differences. When writing schematic vertices such as $(\nabla\eta)(\nabla\phi)(\nabla\phi)$, indices are understood to be contracted in an $\mathrm{SO}(d)$–covariant way.
On the periodic torus we use the backward derivative $(\nabla_{\mu}^{*}f)(n)=(f(n)-f(n-\hat{\mu}))/a$ and the associated backward gradient $(\nabla^{*}F)_{\mu}=\nabla_{\mu}^{*}F$. Then for any scalar $F$, we have
$\sum_{n}(\nabla\!\cdot\eta)(n)\,F(n)\;=\;-\sum_{n}\eta(n)\cdot(\nabla^{*}F)(n),$ (4.33)
with no boundary term on the torus.
All local geometric data entering the field actions are smooth local functions of $\nabla\eta(n)$, more precisely of the frame matrix $E(n)$ built from the $e_{\mu}(n)$. Expanding around the regular embedding therefore yields schematic local expansions
$g_{\mu\nu}(n)=a^{2}\Big(\delta_{\mu\nu}+O(\nabla\eta)\Big),\qquad V(n)=a^{d}\Big(1+O(\nabla\eta)+O(|\nabla\eta|^{2})\Big),$ (4.34)
with coefficients bounded uniformly on $\mathcal{X}_{\mathrm{adm}}$ by admissibility (uniform shape regularity).
Let us consider the scalar $\phi^{4}$ theory for definiteness. Writing the scalar action on a general admissible geometry as $S_{\phi}[x,\phi]=S_{\phi}^{(0)}[\phi]+S^{(1)}_{\phi\eta}[\phi,\eta]+S^{(2)}_{\phi\eta}[\phi,\eta]+\cdots$, we see that the regular-lattice part $S_{\phi}^{(0)}$ is the usual hypercubic action, while the geometry-dependent pieces split into two distinct families already at first order, as follows.
@@
$\begin{split}S^{(1)}_{\phi\eta}&\supset\ \sum_{n}a^{d}\;(\nabla\eta)(n)\,(\nabla\phi)(n)\,(\nabla\phi)(n),\\ S^{(2)}_{\phi\eta}&\supset\ \sum_{n}a^{d}\;|\nabla\eta(n)|^{2}\,(\nabla\phi)^{2}(n),\end{split}$ (4.35)